Factor the equation and use the Quadratic Formula to solve for x. Remember that x must be a real number.

How many unique real values of x solve x4 - 16 = 0?
Answer #1

x4 - 16 = 0

rewrite this as

(x2)2 - 42 = 0

apply difference of two squares formula, we get

(x2 - 4) (x2 + 4)

factoring, we get

(x2 - 4) = x2 - 22

so,

( x + 2) (x - 2)

so,

real factors of x are

x = 2 and x = -2